OpenStudy (babyslapmafro):

Let f(x)=x^2+px+q. Find the values of p and q such that f(2)=2 is an extreme value of f on [0,4]. Is this value a max or min?

OpenStudy (babyslapmafro):

All I have so far is the derivative... f ' (x)= 2x+p

jimthompson5910 (jim_thompson5910):

Hint: f(x)=x^2+px+q is a parabola that opens upward (because the leading coefficient, which is 1, is positive)

OpenStudy (babyslapmafro):

Ok so this function will have an absolute min

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

the min occurs when f '(x) = 0

OpenStudy (babyslapmafro):

When finding the critical value, will the x-value include p. This is my critical value: x=(-p/2)

jimthompson5910 (jim_thompson5910):

or you can say because (2,2) is an extrema, you know that f ' (2) = 0

OpenStudy (babyslapmafro):

oh ok, so p=-4

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so f(x)=x^2+px+q turns into f(x)=x^2-4x+q

OpenStudy (babyslapmafro):

how do you know (2,2) is an extrema?

jimthompson5910 (jim_thompson5910):

it says " ... f(2)=2 is an extreme value of f... "

OpenStudy (babyslapmafro):

oh dang, i wrote typed the problem wrong... f(2)=3 is an extreme value of f

jimthompson5910 (jim_thompson5910):

doesn't matter, things won't change too much

jimthompson5910 (jim_thompson5910):

p is still -4

OpenStudy (babyslapmafro):

ok right because the x value is what is most important for finding p

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you've yet to use the y coordinate of the extrema

OpenStudy (babyslapmafro):

q=7

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (babyslapmafro):

thanks for the help

jimthompson5910 (jim_thompson5910):

So your function is f(x) = x^2 - 4x + 7 You can verify by graphing

jimthompson5910 (jim_thompson5910):

np